OpenStudy (anonymous):

john needs to cut three strips of wood, 11 3/5 inches, 15 1/5 inches, and 9 2/5 inches. He has a strip that is 48 inches long. How long will the leftover piece be after he cuts what he wants?

OpenStudy (anonymous):

11 3/5 15 1/5 9 2/5 ______ 35 6/5 35 +1 + 1/5 = 36 1/5 48 - 36 1/5 = 47 5/5 - 36 1/5 = 11 4/5